# Learning Effective Representations for Retrieval Using Self-Distillation with Adaptive Relevance Margins

### v1

@@ 4. Experimental Setup
Both training and evaluation use the MSMARCO-Passage collection (Nguyen et al., 2016). For training, we use the official set of training triplets, each comprising a query, positive, and negative text passage. For evaluation, we employ relevance annotations from the TREC-DL19 (Craswell et al., 2020b) and TREC-DL20 (Craswell et al., 2020a) datasets, which contain 43 and 54 densely judged queries for the MSMARCO-Passage collection, respectively. Since TREC annotations include graded relevance (ranging from 0 for non-relevant to 3 for perfect), we use the recommended binarization level of (Craswell et al., 2020b) for all binary metrics. Text is tokenized with a maximum sequence length of 30 for queries and 200 for passages (Hofstätter et al., 2021).
@@
We implement our experiment pipeline in PyTorch (Paszke et al., 2019) using the Lightning (Falcon and The PyTorch Lightning team, 2019) and Hugging Face (Wolf et al., 2020) ecosystem. We fine-tune several pretrained BERT-style encoder models, which vary in parameter count, pretraining regime, and embedding kind. The full list is provided in Table 1. Text representations are computed using the [CLS] representation of each document, optionally processed by a linear layer. All models produce 768-dimensional embeddings. We train using the ADAM optimizer with a learning rate of , a weight decay rate of , and batch sizes . Training uses a single Nvidia A100 40GB GPU. To improve training efficiency, we use mixed-precision weights and quantized optimizer states.We provide fully reproducible code and supplementary run files for trained models.333Code: github.com/webis-de/adaptive-relevance-margin-loss
Run files: zenodo.org/records/11197962
@@
We study two evaluation setups: reranking and full ranking. In the reranking setup, which is used for evaluating ablation runs with low overhead, we rerank the top-1000 documents from the official TREC-DL19/20 BM25 baseline runs for each query. For the full ranking setup, replicating real-world use cases, we embed and index the entire MSMARCO-Passage collection using a FAISS (Johnson et al., 2021) HNSW index with optimized 32-byte product quantization, and perform exact re-ranking of the top- results. In both setups, we use cosine similarity between query and document representations for relevance estimation.Similar to prior work, in the full ranking setup we report the measures nDCG@10, representing ranking-oriented tasks, and Recall@1000, representing retrieval-oriented tasks. However, in the reranking setup, the BM25 baseline run artificially restricts the recall obtainable. Therefore, we use Hits@100 as a suitable surrogate for recall that does not rely on all relevant documents being retrievable for each topic.
@@
To contextualize the results of our proposed loss functions, we compare against three different baseline models: (1) BM25 given its nature as standard retrieval baseline and reranking input in our reranked evaluation setting; (2) a BERT-based bi-encoder finetuned with a static margin of (Hofstätter et al., 2021) (3) a BERT-based bi-encoder finetuned with TAS-B batch sampling and teacher margins (Hofstätter et al., 2021; Wang and Zuccon, 2023).
@@ 5. Results & Discussion
We conduct four experiments to study the benefits of using adaptive and distributed margins for fine-tuning bi-encoder retrieval models. First, we conduct an ablation study on static margins to see how adjusting the parameter affects effectiveness, with and without in-batch negatives (Section 5.1). Second, we study the effectiveness of adaptive margins, including ablation studies on batch size and use of in-batch negatives (Section 5.2). Third, we repeat this analysis for distributed margins (Section 5.3). Finally, we compare the optimal parameter settings of each loss variant found in the previous analyses (Section 5.4), and compare them to other state-of-the-art and baseline models (Section 5.5). This comparison provides insights into the overall effectiveness of our proposed approach.
@@ 5.1. Static Targets
In the first ablation experiment, we conduct an exhaustive grid search to find the optimal static target hyperparameter . We fine-tune each of the three pretrained encoders using the static target loss variant, while varying in the range of 0 to 1 in steps of 0.1. To make this exhaustive grid search feasible in terms of computation time, we train each configuration on the same randomly sampled subset of MSMARCO-Passage, containing . Evaluation scores are calculated in the reranking setting. The maximum batch size possible across all models on the available hardware was used (). Experiments with smaller batch size () showed the same trends and relative differences.Figure 3 shows evaluation scores for each combination of and model, with and without using in-batch negatives. Four effects are apparent: (1) Tuning as a hyperparameter yields improvements across all configurations compared to the default choice of . Moreover, the improvement between the default and optimal configuration is distinct, surpassing 0.1 nDCG@10 in almost all cases. Consequently, margin optimization is necessecary to extract maximum effectiveness in each setup. (2) The optimal value of varies depending on the model; each model achieves its highest evaluation score at different points. This indicates that hyperparameter optimization needs to be repeated for each pretrained model, and there is no universal choice for retrieval finetuning as a whole. (3) The optimal margin varies depending on the task; all tested models show a different -value at which the respective evaluation measure, i.e., nDCG@10 or Hits@100, is maximized. (4) Using in-batch negatives does not yield an improvement in effectiveness. For all three tested models and both measures, the effectiveness without in-batch negatives is the same or higher than with in-batch negatives. This suggests that the models do not benefit from the additional training data, likely because it is noisy and provides little information, as negatives in randomly sampled batches are ‘easy’. The effectiveness of the static target loss variant, whether with or without in-batch negatives, is highly susceptible to the training setup used. Factors such as model choice, value of , and measure result in varying levels of effectiveness and exhibit complex interdependencies. At the same time, substantial gains in effectiveness can be achieved through hyperparameter tuning. However, negative mining techniques appear necessary to effectively use in-batch information.
@@ 5.2. Adaptive Targets
In the second ablation experiment, we examine various configurations of the adaptive target approach. Figure 4 shows the evaluation scores computed at every 1000th step during the training process. The training uses the entire MSMARCO-Passage collection. Evaluation scores are calculated in the reranking setting with a batch size of ; similar trends were observed for .
Two key observations can be made: (1) Using adaptive targets is data efficient. The maximum nDCG@10 score is attained after processing approximately half of the data. However, effectiveness declines thereafter, suggesting a divergence from the target task. This issue could potentially be addressed through learning rate scheduling or by terminating training early. (2) The use of in-batch negatives produces mixed results in terms of effectiveness. Regarding nDCG@10, across all configurations, the achieved score is lower when using in-batch negatives, and data efficiency is reduced. The maximum score is reached much later in the training process. However, for Hits@100, only RoBERTa does not benefit from the additional information, and the negative impact on data efficiency is diminished. Adaptive targets can be effectively employed as a hyperparameter-free replacement of the static target variant, yielding comparable evaluation scores. It demonstrates more retrieval-favored characteristics (Hits@100) than the static variant, albeit at the expense of ranking accuracy (nDCG@10).
@@ 5.3. Distributed Targets
In the third ablation experiment, we examine the distributed targets variant. Figure 5 shows evaluation scores calculated in the reranked setting for all three pretrained models. It explicitly includes ablation regarding batch size, i.e., negative sample count.Three key observations can be made for the distributed targets variant: (1) it converges towards the maximum nDCG@10 score even quicker than the adaptive targets variant (see Figure 4). Batch size influences convergence, with smaller batches yielding a more data-efficient result. (2) The use of in-batch information leads to improved results compared to previous in-batch static or adaptive variants. Specifically, nDCG@10 is highest in the distributed setup. (3) Absolute maximum scores are only marginally affected by batch size, with larger batch sizes yielding slightly higher nDCG@10 scores in two out of the three fine-tuned models. Distributed targets also shows to be a robust hyperparameter-free replacement for static targets. They exhibit complementary behavior to adaptive targets, showcasing better ranking accuracy (nDCG@10) while maintaining similar retrieval characteristics (Hits@100).
@@
To contextualize these findings, Figure 6 illustrates the training characteristics of the distributed target variant. It plots the mean, minimum, and maximum values of the similarity matrix throughout the training process. (1) The in-batch target distribution rapidly settles to a mean of 0 (rescaled to 0.5 for loss computation), suggesting that non-relevant documents are, on average, orthogonal in latent space to relevant documents. This indicates an ideal class separation in cosine similarity space. (2) The distribution features outliers in the positive spectrum, i.e., the maximum value (upper line) is further apart from the mean (middle line) than the minimum value (lower line). Consequently, documents highly similar to the relevant one of a query are less common than dissimilar ones, approaching the expected imbalanced ratio of relevant to irrelevant documents for a query. (3) While the initial pretrained similarity distribution, i.e., the first 100 steps, differs significantly among the three pretrained models, their finetuned distributions are largely identical. This suggests that a robust latent space for retrieval can be generalized independently of pretraining regime and architecture. Table 2 presents an exemplary excerpt of a training batch, comprising a query with its associated positive and three negative passages (maximum, minimum, and closest to the mean predicted target). Additionally, the table provides the predicted relevance (), scaled target (), and instance loss () calculated with the fine-tuned DistilBERT model. The theorized loss properties can be observed. The relevance () of documents aligns with their target ordering, i.e., semantic document similarity, showcasing successful self-distillation. Moreover, the less related a negative document is to the positive document, the lower its contribution to the overall loss, showcasing implicit negative mining.
@@
| Wild rice (Z. aquatica), also an annual, grows in the Saint Lawrence River […] | 0.79 | — | — | |
TIvan L. Sander. Northern red oak (Quercus rubra), also known as common […] | 0.34 | 0.69 | 0.06 |
Your answer will appear in the acres field. How to convert acres to ft2 […] | 0.06 | 0.51 | 0.05 |
Colorado Springs Municipal (Colorado Springs, CO) Right now, 14 airlines […] | 0.12 | 0.46 | 0.04 |
@@ 5.4. Comparison of Target Variants
To compare all three variants of the proposed loss function, we train a model for each using parameter choices informed by the previous ablation studies. We use a batch size of and employ exponential learning rate decay (). To prevent the effectiveness degradation observed for adaptive and distributed variants, we implement an early stopping criterion. If no improvement in nDCG@10 score on TREC-DL19 calculated every 500 batches is achieved for 16 consecutive checks training is halted.
| M | T | IB | TREC-DL19 | TREC-DL20 | Time | Data Seen | ||
| nDCG@10 | R@1000 | nDCG@10 | R@1000 | (hh:mm) | (# Triples) | |||
@@
| Loss Variant | B. Sampl. | In-B. | Source | TREC-DL19 | TREC-DL20 | Data Seen | |||
| nDCG@10 | Recall@1000 | nDCG@10 | Recall@1000 | (# Triples) | |||||
| Self-D. | Adaptive | Rand. | ✗ | Ours | 4.93M | ||||
| Adaptive | Rand. | ✓ | Ours | 1.98M | |||||
| Distributed | Rand. | ✓ | Ours | 7.30M | |||||
| Teacher Distillation | MarginMSE | Rand. | ✗ | Hofstätter et al. (2021) | 22.4M | ||||
| Rand. | ✓ | Hofstätter et al. (2021) | 22.4M | ||||||
| TAS-B | ✗ | Hofstätter et al. (2021) | 22.4M | ||||||
| TAS-B | ✓ | Hofstätter et al. (2021) | 22.4M | ||||||
| Rand. | ✗ | Wang and Zuccon (2023) | 14.6M | ||||||
| Rand. | ✓ | Wang and Zuccon (2023) | 14.6M | ||||||
| TAS-B | ✗ | Wang and Zuccon (2023) | 14.6M | ||||||
| TAS-B | ✓ | Wang and Zuccon (2023) | 14.6M | ||||||
@@
| Static (Tuned) | Rand. | ✗ | Ours | 5.2M | |||||
| Static (Tuned) | Rand. | ✓ | Ours | 2.2M | |||||
| Static () | Rand. | ✗ | Hofstätter et al. (2021) | 22.4M | |||||
Table 3 lists the obtained evaluation scores in the full ranking setup on both TREC-DL19 and TREC-DL20 for each of the three pretrained models investigated, each of the three loss variants, and (for static and adaptive targets) with and without in-batch negatives. Statistical equivalence was tested among all approaches per model to determine if the difference of each approach to the respective best per measure and model is meaningful (paired TOST, , , [TREC-DL19], [TREC-DL20], Bonferroni correction applied per test group). Equivalence of approaches is indicated in Table 3 using a symbol. Additionally, the table lists the number of training samples seen and the total training time per approach to compare their data/compute efficiency. For all static target variants, this includes compute time spent on hyperparameter search to identify the optimal target value.For nDCG@10, the static target approach without in-batch negatives achieves the highest scores across all models and both datasets. However, the distributed targets approach, while scoring slightly lower in all setups is still statistically equivalent, except on TREC-DL20 with RoBERTa as the base model. For Recall@1000, the adaptive target variant with in-batch negatives achieves the highest scores across all models and both datasets, except for distilBERT on TREC-DL19, where the static approach without in-batch negatives scores higher. Additionally, in half of the test cases (distilBERT, RoBERTa on TREC-DL19), the distributed targets are statistically equivalent to the best score. Therefore, adaptive and distributed targets are suitable hyperparameter-free replacements for tuned static targets for retrieval (Recall@1000) and ranking tasks (nDCG@10).These results also highlight a large difference in training efficiency. Since the static target setup requires hyperparameter tuning, its total training time is between 3x (distilBERT) and 6x (MPNet) higher than the distributed target approach, and between 7x and 15x higher than the adaptive target approach. When comparing efficiency in terms of training triplets seen, two further trends are identified: using in-batch information increases data efficiency, and self-distillation approaches are more data efficient than their static counterparts in all but one case (distilBERT). In-batch information is also detrimental for static variants, suggesting that our approach to implicit negative mining improves effectiveness. The observed gains in compute- and data-efficiency strengthen the applicability of our proposed approach, as it is not only on par in terms of effectiveness but also does not require large amounts of data or compute. Additionally, the choice of pretrained model does not yield meaningful differences in training outcome; across all evaluated setups, the absolute scores show no clear difference between the three pretrained models, and the relative trends concerning the individual loss variants are also stable across all pretrained models.
5.5. Comparison to Baselines

### v2

@@ 4. Experimental Setup
We employ the MSMARCO-Passage collection (Nguyen et al., 2016). For training, we use the official training set of training triplets, each comprising a query, positive, and negative document (i.e., text passage in this case). For evaluation, we employ relevance annotations from the TREC-DL 19 (Craswell et al., 2020b) and TREC-DL 20 (Craswell et al., 2020a) datasets, which contain 43 and 54 densely judged queries for the MSMARCO-Passage collection, respectively. Since TREC annotations include graded relevance (ranging from 0 for non-relevant to 3 for maximum relevance), we use the recommended binarization level of (Craswell et al., 2020b) for all binary metrics. Text is tokenized with a maximum sequence length of 30 for queries and 200 for passages (Hofstätter et al., 2021).
@@
We implement our experiment pipeline in PyTorch (Paszke et al., 2019) using the Lightning (Falcon and The PyTorch Lightning team, 2019) and Hugging Face (Wolf et al., 2020) libraries. We fine-tune several pre-trained BERT-style encoder models, which vary in parameter count, pre-training regime, and embedding kind. The full list is provided in Table 1. Text representations are computed using the [CLS] representation of each document, optionally processed by a linear layer. All models produce 768-dimensional embeddings. We train using the ADAM optimizer with a learning rate of , a weight decay rate of , and batch sizes .
Hardware.
Training is performed on a single Nvidia A100 40GB GPU, which renders our training setup attainable for most academic and small-scale industrial research environments. To improve training efficiency, we use mixed-precision weights and quantized optimizer states, reducing memory footprint by approximately 50% compared to full-precision training while maintaining comparable model effectiveness.
@@
We study two evaluation setups: re-ranking and full ranking. In the re-ranking setup, which is used for evaluating runs with low overhead, we re-rank the top-1000 documents from the official TREC-DL 19/20 BM25 baseline runs for each query. For the full ranking setup, replicating real-world use cases, we embed and index the entire MSMARCO-Passage collection using a FAISS (Johnson et al., 2021) HNSW index with optimized 32-byte product quantization, and perform exact re-ranking of the top- results. In both setups, we use cosine similarity between query and document representations for relevance estimation.Similar to prior work, in the full ranking setup we report the measures nDCG@10, representing ranking-oriented tasks, and Recall@1000, representing retrieval-oriented tasks. However, in the re-ranking setup, the BM25 baseline run artificially restricts the recall obtainable. Therefore, we use Hits@100 as a suitable surrogate for recall that does not rely on all relevant documents being retrievable for each topic.
@@
To contextualize the results of our proposed loss functions, we compare against three different baseline models: (1) BM25 given its nature as standard retrieval baseline and re-ranking input in our re-ranked evaluation setting; (2) a BERT-based bi-encoder fine-tuned with a static margin of (Hofstätter et al., 2021) (3) a BERT-based bi-encoder fine-tuned with TAS-B batch sampling and teacher margins (Hofstätter et al., 2021; Wang and Zuccon, 2023).
@@ 5. Results & Discussion
We conduct four experiments to study the benefits of using adaptive and distributed margins for fine-tuning bi-encoder retrieval models. First, we conduct an experiment on static margins to see how adjusting the parameter affects effectiveness, with and without in-batch negatives (Section 5.1). Second, we study the effectiveness of adaptive margins, including experiments on batch size and ablating in-batch negatives (Section 5.2). Third, we repeat this experiment for distributed margins (Section 5.3). Finally, we compare the optimal parameter settings of each loss variant found in the previous experiments (Section 5.4), and compare them to other state-of-the-art and baseline models (Section 5.5). This comparison provides insights into the overall effectiveness of our proposed approach.
@@ 5.1. Static Targets
In the first experiment, we conduct a grid search to find the optimal static target hyperparameter . We fine-tune each of the three pre-trained encoders using the static target loss variant, while varying in the range of 0 to 1 in steps of 0.1. To make this grid search feasible in terms of computation time, we train each configuration on the same randomly sampled subset of MSMARCO-Passage, containing . Evaluation scores are calculated in the re-ranking setting. The maximum batch size possible across all models on the available hardware was used (). Experiments with smaller batch size () showed the same trends and relative differences.
Figure 3 shows evaluation scores for each combination of and model, with and without using in-batch negatives. Four effects are apparent: (1) Tuning as a hyperparameter yields improvements across all configurations compared to the default choice of . Moreover, the improvement between the default and optimal configuration is distinct, surpassing 0.1 nDCG@10 in almost all cases. Consequently, margin optimization is necessary to extract maximum effectiveness in each setup. (2) The optimal value of varies depending on the model; each model achieves its highest evaluation score at different points. This indicates that hyperparameter optimization needs to be repeated for each pre-trained model, and there is no universal choice for retrieval fine-tuning as a whole. (3) The optimal margin varies depending on the task; all tested models show a different -value at which the respective evaluation measure, i.e., nDCG@10 or Hits@100, is maximized. (4) Using in-batch negatives does not yield an improvement in effectiveness. For all three tested models and both measures, the effectiveness without in-batch negatives is the same or higher than with in-batch negatives. This suggests that the models do not benefit from the additional training data, likely because it is noisy and provides little information, as negatives in randomly sampled batches are ‘easy’. The effectiveness of the static target loss variant, whether with or without in-batch negatives, is highly susceptible to the training setup used. Factors such as model choice, value of result in varying and interdependent levels of effectiveness. Substantial gains in effectiveness can be achieved through hyperparameter tuning. However, negative mining techniques beyond simply random sampling appear necessary to effectively use in-batch information.
@@ 5.2. Adaptive Targets
In the second experiment, we examine various configurations of the adaptive target approach. Figure 4 shows the evaluation scores computed at every 1000th step during the training process. The training uses the entire MSMARCO-Passage collection. Evaluation scores are calculated in the re-ranking setting with a batch size of ; similar trends were observed for .Two key observations can be made: (1) Using adaptive targets is data efficient. The maximum nDCG@10 score is attained after processing approximately half of the data. However, effectiveness declines thereafter, suggesting a divergence from the target task. (2) The use of in-batch negatives produces mixed results in terms of effectiveness. Regarding nDCG@10, across all configurations, the achieved score is lower when using in-batch negatives, and data efficiency is reduced. The maximum score is reached much later in the training process. However, for Hits@100, only RoBERTa does not benefit from the additional information, and the negative impact on data efficiency is diminished. Adaptive targets can be effectively employed as a hyperparameter-free replacement of the static target variant, yielding comparable evaluation scores. It demonstrates more retrieval-favored characteristics (Hits@100) than the static variant, albeit at the expense of ranking effectiveness (nDCG@10).
@@ 5.3. Distributed Targets
In the third experiment, we examine the distributed targets variant. Figure 5 shows evaluation scores calculated in the re-ranked setting for all three pre-trained models. It includes a variation of the batch size, i.e., negative sample count.Three key observations can be made for the distributed targets variant: (1) it converges towards the maximum nDCG@10 score even quicker than the adaptive targets variant (see Figure 4). Batch size influences convergence, with smaller batches yielding a more data-efficient result. (2) The use of in-batch information leads to improved results compared to previous in-batch static or adaptive variants. Specifically, nDCG@10 is highest in the distributed setup. (3) Absolute maximum scores are only marginally affected by batch size, with larger batch sizes yielding slightly higher nDCG@10 scores in two out of the three fine-tuned models. Distributed targets also shows to be a robust hyperparameter-free replacement for static targets. They exhibit complementary behavior to adaptive targets, showcasing better ranking accuracy (nDCG@10) while maintaining similar retrieval characteristics (Hits@100).
@@
To contextualize these findings, Figure 6 illustrates the training characteristics of the distributed target variant. It plots the mean, minimum, and maximum values of the similarity matrix throughout the training process. Four observations stand out: (1) The in-batch target distribution rapidly settles to a mean of 0 (rescaled to 0.5 for loss computation), suggesting that non-relevant documents are, on average, orthogonal in latent space to relevant documents. This indicates an ideal class separation in cosine similarity space. (2) The distribution features outliers in the positive spectrum, i.e., the maximum value (upper line) is further apart from the mean (middle line) than the minimum value (lower line). Consequently, documents highly similar to the relevant one of a query are less common than dissimilar ones, approaching the expected imbalanced ratio of relevant to irrelevant documents for a query. (3) While the initial pre-trained similarity distribution, i.e., the first 100 steps, differs significantly among the three pre-trained models, their fine-tuned distributions are largely identical. This suggests that a robust latent space for retrieval can be generalized independently of pre-training regime and architecture. Table 2 presents an exemplary excerpt of a training batch, comprising a query with its associated positive and three negative passages (maximum, minimum, and closest to the mean predicted target). Additionally, the table provides the predicted relevance , scaled target , and instance loss calculated with the fine-tuned distilBERT model. The theorized loss properties can be observed. The relevance of documents aligns with their target ordering, i.e., semantic document similarity, showcasing successful self-distillation. Moreover, the less related a negative document is to the positive document, the lower its contribution to the overall loss, showcasing implicit negative mining.
@@
| Wild rice (Z. aquatica), also an annual, grows in the Saint Lawrence River, the state of Florida, and on the Atlantic and Gulf coasts of the United States. Texas wild rice is a perennial plant found only in a small area along the […] | 0.79 | — | — | |
| Ivan L. Sander. Northern red oak (Quercus rubra), also known as common red oak, eastern red oak, mountain red oak, and gray oak, is widespread in the East on a variety of soils and topography, often forming pure stands. […] | 0.34 | 0.69 | 0.06 | |
| Your answer will appear in the acres field. How to convert acres to ft2 : Use the conversion calculator titled Convert acres to ft2. Enter a value in the acres field and click on the Calculate square feet button. Your answer will […] | 0.06 | 0.51 | 0.05 | |
| 1 Colorado Springs Municipal (Colorado Springs, CO) 2 Right now, 14 airlines operate out of Colorado Springs Municipal. 3 Colorado Springs Municipal offers nonstop flights to 11 cities. Every week, at least 441 domestic flights and 7 international flights depart […] | 0.12 | 0.46 | 0.04 |
@@ 5.4. Comparison of Target Variants
To compare all three variants of the proposed loss function, we train a model for each using parameter choices informed by the previous experiments. We use a batch size of and employ exponential learning rate decay . To prevent the effectiveness degradation observed for adaptive and distributed variants, we implement an early stopping criterion. If no improvement in nDCG@10 score on TREC-DL 19 calculated every 500 batches is achieved for 16 consecutive checks, training is halted.
| M | T | IB | nDCG@10 | Recall@1000 | Time | Data Seen | ||
| DL19 | DL20 | DL19 | DL20 | (hh:mm) | (# Triplets) | |||
@@
Table 3 lists the obtained evaluation scores in the full ranking setup on both TREC-DL 19 and TREC-DL 20 for each of the three pre-trained models investigated, each of the three loss variants, and (for static and adaptive targets) with and without in-batch negatives. Statistical equivalence was tested among all approaches per model to determine if the difference of each approach to the respective best per measure and model is meaningful (paired TOST, , , [TREC-DL 19], [TREC-DL 20], Bonferroni correction applied per test group). Equivalence of approaches is indicated in Table 3 using a symbol. Additionally, the table lists the number of training samples seen and the total training time per approach to compare their data/compute efficiency. For all static target variants, this includes compute time spent on hyperparameter search to identify a reasonable target value. Note that TREC-DL 19 is used both to tune hyperparameter values, and to report model effectiveness with. This is deliberate, in order to illustrate a reasonable upper bound on the effectiveness of hyperparameter-based approaches under train/test leakage conditions. TREC-DL 20 presents scores in a neutral setting.For nDCG@10, the static target approach without in-batch negatives achieves the highest scores across all models and both datasets. However, the distributed targets approach, while scoring slightly lower in all setups is still statistically equivalent, except on TREC-DL 20 with RoBERTa as the base model. This is especially interesting, since the static target approach was privy to test data during hyperparameter optimization, which the distributed approach does not need. For Recall@1000, the adaptive target variant with in-batch negatives achieves the highest scores across all models and both datasets, except for distilBERT on TREC-DL 19, where the static approach without in-batch negatives scores higher. Additionally, in half of the test cases (distilBERT, RoBERTa on TREC-DL 19), the distributed targets are statistically equivalent to the highest score. Therefore, adaptive and distributed targets are suitable hyperparameter-free replacements for tuned static targets for retrieval (Recall@1000) and ranking tasks (nDCG@10).These results also highlight a large difference in training efficiency. Since the static target setup requires hyperparameter tuning, its total training time is between 3x (distilBERT) and 6x (MPNet) higher than the distributed target approach, and between 7x and 15x higher than the adaptive target approach. When comparing efficiency in terms of training triplets seen, two further trends are identified: using in-batch information increases data efficiency, and self-distillation approaches are more data efficient than their static counterparts in all but one case (distilBERT). In-batch information is also detrimental for static variants, suggesting that our approach to implicit negative mining improves effectiveness. The observed gains in compute- and data-efficiency strengthen the applicability of our proposed approach, as it is not only on par in terms of effectiveness but also does not require large amounts of data or computing time. The inclusion of an additional doc-to-doc similarity computation, making available for training signal in each backpropagation step, could serve as explanation for both the data efficiency, as well as the tendency to overfit in adaptive and distributed margins. The choice of pre-trained model does not yield meaningful differences in training outcome; across all evaluated setups, the absolute scores show no clear difference between the three pre-trained models, and the relative trends concerning the individual loss variants are also stable across all pre-trained models.
| Loss | B. Smpl. | IB | Src. | nDCG@10 | Recall@1000 | Data | |||
| DL19 | DL20 | DL19 | DL20 | (# Tripl.) | |||||
| Self-D. | A | Rand. | ✗ | Ours | 4.93M | ||||
| A | Rand. | ✓ | Ours | 1.98M | |||||
| D | Rand. | ✓ | Ours | 7.30M | |||||
| Teacher Distillation | MarginMSE | Rand. | ✗ | 22.4M | |||||
| Rand. | ✓ | (Hofstätter et al., 2021) | 22.4M | ||||||
| TAS-B | ✗ | (Orig.) | 22.4M | ||||||
| TAS-B | ✓ | 22.4M | |||||||
| Rand. | ✗ | 14.6M | |||||||
| Rand. | ✓ | (Wang and Zuccon, 2023) | 14.6M | ||||||
[…]
| TAS-B | ✓ | 14.6M | |||||||
@@
| S (T) | Rand. | ✗ | Ours | 5.2M | |||||
| S (T) | Rand. | ✓ | Ours | 2.2M | |||||
| S () | Rand. | ✗ | (Hofstätter et al., 2021) | 22.4M | |||||
| S (T) denotes tuned, S () denotes untuned static margins. | |||||||||
5.5. Comparison to TAS-B